OpenStudy (anonymous):

At Trent's garage, 3 cases of brake fluid last 4 weeks. Trent has 5 cases. How much more brake fluid should Trent order so that he has enough for 13 weeks? Round your answer up to the nearest whole number.

OpenStudy (anonymous):

3/4 = (x + 5)/13 39 = 4x + 20 19 = 4x x ≈ 5